# Impact of Exchange-Correlation Functionals on Predictions of Phonon Hydrodynamics: A Study of Fluorides, Chlorides, and Hydrides

### v2

@@ Abstract
We employ density functional theory calculations to examine the effect of various exchange-correlation (XC) functionals, including the Perdew-Burke-Ernzerhof generalized gradient approximation (PBE), the modified Perdew-Burke-Ernzerhof generalized gradient approximation (PBESOL), and the local density approximation (LDA), on the electrical, mechanical, and thermal properties of sodium fluoride (NaF), lithium fluoride (LiF), potassium fluoride (KF), sodium chloride (NaCl), potassium chloride (KCl), lithium hydride (LiH), sodium hydride (NaH), and potassium hydride (KH). The lattice thermal conductivity is computed based on an iterative solution of the Boltzmann transport equation (BTE). Based on Guyer’s criterion and direct solutions to the linearized BTE, we determine the ballistic, phonon hydrodynamics and diffusive regimes as function of length scale and temperature. In addition to confirming previous predictions of phonon hydrodynamics in NaF and LiF, we report novel predictions of phonon hydrodynamics in NaH, LiH, KH, KF, NaCl and KCl. The impact of isotopes on the calculated lattice thermal conductivity and phonon hydrodynamics windows is also reported. The impact of Meta-GGA and hybrid functionals is also discussed. We find that the selection of a functional impacts the prediction of thermal conductivity and the window for observation of phonon hydrodynamics.
keywords:
Exchange-correlation functionals , Density functional theory , Phonon hydrodynamics , Second sound , Guyer’s criterion , Boltzmann transport equation (BTE)
journal: Computational Materials Science
[…]
@@ 1 Introduction
Based on the original insights of Hohenberg and Kohn, who first proved that the ground-state electron density uniquely determines the energy of a many-electron system [kohn1965self], density functional theory (DFT) has emerged as a standard tool in computational materials science. DFT’s success can be attributed to its ability to balance accuracy and computational cost [van2014density]. At the heart of this balance lies the exchange correlation (XC) functional, which approximates many-body electron interactions. This functional operates on a foundation that facilitates a ladder of approximations. The choice of XC approximation is critical, as it largely determines the trade-off between accuracy and computational expense in DFT simulations [dick2020machine].
@@
The bottom rung in the ladder of functional approximations is the local density approximation (LDA), which approximates the exchange correlation energy based on a uniform electron gas model [perdew1996generalized]. Advancements led to the formulation of generalized gradient approximations (GGA), which extend the LDA approach by incorporating gradient corrections to the electron density, enhancing the accuracy of DFT calculations. Several GGA functions have since been developed since there are multiple ways to characterize the gradient of electron density [perdew1996generalized, chen2006b3lyp, perdew2008restoring]. The two most common GGA functionals are the Perdew-Burke-Ernzerhof generalized gradient approximation (PBE) [perdew1996generalized] and the modified Perdew-Burke-Ernzerhof generalized gradient approximation (PBESOL) [perdew2008restoring]. PBE includes the gradient of the electron density in the exchange correlation function and improves upon LDA by capturing some non-local effects, but there remain limitations, particularly for systems with strongly localized electrons (e.g., transition metal complexes [zhang2018performance]). The third most common functional is meta-generalized gradient approximation (meta-GGA), which augments the functional dependence by including higher-order terms like the second derivative of the electron density [marques2003primer]. In general, meta-GGAs (e.g., SCAN) offer improvements over GGAs in terms of accuracy for a variety of properties, including thermochemistry, and provide reliable predictions of structures and energies for a wide range of molecular and material bonding types [sun2016accurate]. Another noteworthy functional is the HSE (Heyd-Scuseria-Ernzerhof) hybrid functional which combines the traditional DFT exchange-correlation functional with a fraction of Hartree-Fock (HF) exchange [heyd2003hybrid] via a mixing parameter that determines the contribution of HF exchange to the total exchange-correlation energy. However, this functional requires greater computational resources than the simple PBE functional.
@@
Given these approximations, deviations between experimental and theoretical predictions are expected. It is well known that the LDA under-predicts the bandgap in silicon by about 50 % [persson2006improved, ishii2010all, zhao2019prediction]. However, even with this significant discrepancy, accurate DFT-based predictions of the thermal conductivity of silicon have been obtained [lindsay2013first, parks2020uncertainty, mortazavi2021accelerating]. For instance, Jain and McGaughey showed that the choice of pseudopotential and XC functional in thermal conductivity predictions can impact accuracy by up to 17% [jain2015effect].
[…]
@@ 2 Theory
Based on approximate solutions to the phonon Boltzmann Transport Equation (BTE), a condition for the existence of phonon hydrodynamics was proposed by Guyer and Krumhansl, henceforth referred to as Guyer’s criterion [guyer1966thermal]. This condition states that phonon hydrodynamics can be observed if the average boundary scattering rate $\langle\tau_{B}^{-1}\rangle$ is greater than the average Umklapp scattering rate $\langle\tau_{U}^{-1}\rangle$ but less than the average Normal scattering rate $\langle\tau_{N}^{-1}\rangle$ at a given temperature and for a characteristic length [guyer1966thermal]:
$\langle\tau_{N}^{-1}\rangle\;>\;\langle\tau_{B}^{-1}\rangle\;>\;\langle\tau_{U}^{-1}\rangle$ (1)
In other words, phonon hydrodynamics can be observed if Normal scattering dominates over boundary scattering and Umklapp scattering. The average resistive scattering rate $\langle\tau_{R}^{-1}\rangle$ (sum of Umklapp and isotope scattering rates) can also be used instead of the average Umklapp scattering rate:
$\langle\tau_{N}^{-1}\rangle\;>\;\langle\tau_{B}^{-1}\rangle\;>\;\langle\tau_{R}^{-1}\rangle$ (2)
@@
$\bigl\langle\tau^{-1}\bigr\rangle=\frac{\sum_{i=1}^{N}\tau_{i}^{-1}\,C_{i}}{\sum_{i=1}^{N}C_{i}}=\frac{\sum_{i=1}^{N}\tau_{i}^{-1}\,C_{i}}{C}$ (3)
where $C_{i}$ and $C$ are the heat capacity for a given phonon mode $i$ and the total heat capacity of all phonon modes and $N$ is the number of q-points in the discretized Brillouin zone, respectively. Other expressions for average scattering rates may be selected, but here we follow previous studies [cepellotti2015phonon, ghosh2020phonon, ghosh2022phonon]. Details of the DFT and BTE calculations can be found in Appendix A of the supplementary material.
@@ 3.1 General Properties
Fluoride, chloride and hydride compounds crystallize in the rock-salt structure (face-centered cubic, space group $Fm\overline{3}m$) [hoffmann2020introduction]. Each cation and anion is octahedrally coordinated by six neighbors. NaF, LiF, KF, KCl, LiH, NaH, and KH all adopt the NaCl-type lattice with similar octahedral frameworks, as shown in Fig.  1. A comprehensive analysis, given in Table 1, of the structural, dielectric, vibrational, and electronic properties highlights the relative performance of the PBE, PBESOL, and LDA functionals. The calculated results are systematically compared against available experimental and theoretical data. For fluorides and chlorides, PBE slightly overestimates and LDA underestimates the lattice constant, while PBESOL is the closest to experiments, often deviating by less than 1% from experimental values. For instance, PBESOL lattice constants for KF (5.31 Å) and KCl (6.24 Å) are in excellent agreement with their experimental measurement results of 5.31 Å and 6.25 Å [pies2006crystal]. Conversely, for the hydrides, the PBE functional tends to yield better results, as seen with LiH where the PBE value (4.01 Å) is the closest to the experimental value of 4.083 Å [pretzel1960properties]. The lattice constant increases monotonically from Li to Na to K for each anion, corresponding to the increasing ionic radii.
@@
The dielectric-constant tensor and Born effective charges are calculated based on the density functional perturbation theory (DFPT) to capture the longitudinal and transverse optical phonon (LO-TO) splitting at the gamma point [liang2018lattice]. For fluorides, chlorides and hydrides (NaF, LiF, LiH, etc.), the underlying Bravais lattice is cubic; hence, there is only one independent component in the Born effective charge tensor. In all cases, Born effective charges are nearly ionic ($\lvert Z^{\ast}\rvert\approx 1$), exemplified by NaCl (1.11, 1.09, 1.071.11,\,1.09,\,1.07 vs. $1.099$ [shukla2000ab]) and LiH (1.021.021.031.03 vs. $1.11$ [blat1991calculations]). For non-magnetic, transparent solids, a larger electronic polarizability leads to a larger refractive index ($n$) according to the Lorentz–Lorenz relation: $(n^{2}-1)/(n^{2}+2)=N\alpha/(3\varepsilon_{0})$, and in the optical limit with $\mu_{r}\approx 1$ the high-frequency dielectric constant satisfies $\varepsilon_{\infty}\approx n^{2}$ [griffiths2023introduction]. Across anions at fixed cation, the electronic polarizability increases in the sequence $\mathrm{F^{-}}<\mathrm{Cl^{-}}<\mathrm{H^{-}}$, so the high-frequency dielectric $\varepsilon_{\infty}$ constant and thus the refractive index ($n\approx\sqrt{\varepsilon_{\infty}}$ ) rises accordingly. The calculated high-frequency dielectric constants follow chemical intuition: fluorides are the least polarizable ($\varepsilon_{\infty}$ = 1.81.82.12.1), chlorides are higher ($\varepsilon_{\infty}$ = 2.32.32.72.7), and hydrides are the largest ($\varepsilon_{\infty}$ = 2.72.74.84.8).
@@
A consistent weakness remains in the prediction of the electronic band gap ($E_{g}$) using these functionals. For example, the calculated values for KF ($\sim 6\,\mathrm{eV}$) and KCl ($\sim 5\,\mathrm{eV}$) are significantly lower than their respective experimental measurements of $10.9\,\mathrm{eV}$ and $8.7\,\mathrm{eV}$ [pantelides1975electronic]. It was previously found that other functionals, such as the modified Becke–Johnson (mBJ) exchange potential, can be used to obtain a more accurate estimate of the band gap [tran2009accurate, camargo2012performance, messaoudi2015band]. The electronic band structures of fluoride, chloride and hydride materials can be found in Appendix B of the supplementary material, where the Fermi energy ($E_{f}$) is shifted to zero eV.
@@
The dispersion curves for the eight fluoride, chloride and hydrides, seen in Fig.  2, are qualitatively consistent across PBE, PBESOL, and LDA and align with the experimental measurement results. The phonon frequency values of LO and TO modes at the gamma point are also given in Table 1. LDA leads to slightly higher phonon frequencies compared to PBE and PBESOL. Nevertheless, the predictions from all three functionals are in good agreement with experimental measurements along high-symmetry lines. A clear hierarchy emerges from the band shapes. Within each anion family, replacing Li with Na and then K progressively softens both acoustic and optical branches and lowers sound velocities. Across a fixed cation, moving from chloride to fluoride to hydride pushes the optical manifold upward and opens a pronounced acoustic–optical gap. The hydrides (LiH, NaH, KH) show the widest gap and the flattest high-frequency optic branches, reflecting the large mass contrast and strong long-range Coulomb coupling [toberer2011phonon, driscoll2022long]. The large LO-TO splitting frequency values are due to strong polar phonon frequencies (creation of an oscillating electric field inside a crystal caused by phonons due to local charge polarization) [bandyopadhyay2012physics]. The $\Gamma$-point LO–TO separations quantify these patterns. Using the tabulated PBE values, the hydrides have the largest splittings, with $\Delta_{\text{LO--TO}}$ of $14.87\,\mathrm{THz}$ for LiH, $11.78\,\mathrm{THz}$ for NaH, and $9.85\,\mathrm{THz}$ for KH. Fluorides are intermediate, with $10.54\,\mathrm{THz}$ for LiF, $4.89\,\mathrm{THz}$ for NaF, and $4.07\,\mathrm{THz}$ for KF. Chlorides are the smallest, with $2.87\,\mathrm{THz}$ for NaCl and $2.12\,\mathrm{THz}$ for KCl. The PDOS curves reflect these trends.
@@
Figure 2: Calculated phonon dispersion and phonon density of states for: (a) NaF, black circles are the measured data from Ref. [buyers1967lattice]. (b) LiF, black circles are the measured data from Ref. [dolling1968lattice]. (c) KF, black circles are the measured data from Ref. [buhrer1970lattice]. (d) NaCl, black circles are the measured data from Ref. [schmunk1970lattice]. (e) KCl, black circles are the measured data from Ref. [copley1969lattice]. (f) LiH, black circles are the measured data from Ref. [verble1968lattice]. (g) NaH. (h) KH. Calculations were done with PBE (black), PBESOL (red), and LDA (blue).
@@
Table 1: Lattice constant $a$ (Å), high-frequency dielectric constant $\varepsilon_{\infty}$, Born effective charge $Z^{*}$, phonon frequencies at $\Gamma$ $\omega$ (THz), and band gap $E_{g}$ (eV) of calculated (PBE, PBESOL, LDA), other theoretical data, and experimental data for NaF, LiF, KF, NaCl, KCl, LiH, NaH and KH.
@@
$Z^{*}$ 1.02 1.01 0.99 1.019e, 0.956i 1.02j
$\omega_{\mathrm{LO}}$ 11.67 12.01 12.81 11.90e 12.65k
$\omega_{\mathrm{TO}}$ 6.78 7.24 8.32 7.7c, 6.95e 7.86h, 7.51k
$E_{g}$ 6.02 6.11 6.34 12.0a, 6.12 (GGA)e, 7.11 (EV)e, 11.69 (mBJ)e, 5.94l 11.50m
@@
$Z^{*}$ 1.05 1.05 1.04 0.998i, 1.03o, 1.06p
$\omega_{\mathrm{LO}}$ 18.63 19.18 20.09 19.89o, 19.92p 19.70t
$\omega_{\mathrm{TO}}$ 8.09 9.06 10.28 9.0c, 10.61o, 10.29p 9.54h, 9.15t
$E_{g}$ 8.82 9.04 9.64 9.11l, 8.727p, 7.64 (LDA)u, 9.15 (PBE)u, 8.7v 13.6w, 14.2x
@@
$\omega_{\mathrm{LO}}$ 9.05 9.45 10.18 9.55aa 9.55y
$\omega_{\mathrm{TO}}$ 4.98 5.42 6.33 5.8c, 5.89aa 6.05h, 5.57y
$E_{g}$ 5.96 6.15 6.36 6.14l 10.9ab
@@
$a$ 5.70 5.61 5.47 5.80c, 5.70d, 5.701 (FP -LAPW)e, 5.695 (PP)e, 5.52ac, 5.698ad, 5.471 (LDA)ae, 5.698 (GGA)ae 5.57c, 5.60f, 5.60ac, 5.595ae
$\varepsilon_{\infty}$ 2.48 2.57 2.67 2.47e, 2.546af 2.35h
$Z^{*}$ 1.11 1.09 1.07 1.105e, 0.985i, 1.096af 1.099i
$\omega_{\mathrm{LO}}$ 7.34 7.60 8.07 7.46e 7.93ag
$\omega_{\mathrm{TO}}$ 4.47 4.93 5.63 4.7c, 4.62e 5.34h, 5.17ag
$E_{g}$ 4.98 4.99 5.04 5.00l, 5.00 (GGA)e, 6.07 (EV)e, 8.41 (mBJ)e 8.75ah, 9.0ai
@@
$a$ 6.38 6.24 6.07 6.57c, 6.40d, 6.30ac 6.25f, 6.24ac
$\varepsilon_{\infty}$ 2.28 2.38 2.51 2.365af 2.20h, 2.15aj
$Z^{*}$ 1.13 1.12 1.11 1.129af
$\omega_{\mathrm{LO}}$ 5.89 6.16 6.66 6.34aj
$\omega_{\mathrm{TO}}$ 3.77 4.15 4.85 4.1c, 4.41aj 4.53h, 4.36aj
$E_{g}$ 5.05 5.11 5.12 5.32l, 5.5ak 8.7ab
@@
$Z^{*}$ 1.02 1.02 1.03 1.046i, 1.020am, 1.03ar 1.11ar, 0.991as
$\omega_{\mathrm{LO}}$ 32.43 32.37 32.96 28.3ar 32.21an, 33.50ar, 33.60as, 33.58at
$\omega_{\mathrm{TO}}$ 17.56 18.52 19.41 18.1ar, 17.86au 18.14an, 18.40ar, 17.76as, 17.70at
$E_{g}$ 2.98 2.73 2.54 3.00 (GGA)al, 2.61 (LDA)al, 4.75 (GW)al, 2.53au 5.00av, 4.92aw
Table 1 (continued)
Property PBE PBESOL LDA Other theoretical data Experiments
@@
$Z^{*}$ 0.97 0.96 0.95 0.97am, 0.96ax, 0.990 (GGA)ba, 0.974 (LDA)ba
$\omega_{\mathrm{LO}}$ 26.75 26.65 27.52 27.06am
$\omega_{\mathrm{TO}}$ 14.97 15.44 16.83 16.6 (CN)az, 13.9 (QH)az 15az
$E_{g}$ 3.76 3.61 3.55 3.79 (GGA)al, 3.42 (LDA)al, 5.87 (GW)al, 4.90ax, 5.68 (GW)bc, 3.39 (LDA)bc

### v3

@@ Abstract
We employ density functional theory calculations to examine the effect of various exchange-correlation (XC) functionals, including the Perdew-Burke-Ernzerhof generalized gradient approximation (PBE), the modified Perdew-Burke-Ernzerhof generalized gradient approximation (PBEsol), and the local density approximation (LDA), on the electrical, mechanical, and thermal properties of sodium fluoride (NaF), lithium fluoride (LiF), potassium fluoride (KF), sodium chloride (NaCl), potassium chloride (KCl), lithium hydride (LiH), sodium hydride (NaH), and potassium hydride (KH). The lattice thermal conductivity is computed based on an iterative solution of the Boltzmann transport equation (BTE). Based on Guyer’s criterion and direct solutions to the linearized BTE, we determine the ballistic, phonon hydrodynamics, and diffusive regimes as a function of length scale and temperature. In addition to confirming previous predictions of phonon hydrodynamics in NaF and LiF, we report novel predictions of phonon hydrodynamics in NaH, LiH, KH, KF, NaCl, and KCl. The impact of isotopes on the calculated lattice thermal conductivity and phonon hydrodynamics windows is also reported. The impact of Meta-GGA and hybrid functionals is also discussed. We find that the selection of a functional impacts the prediction of thermal conductivity and the window for observation of phonon hydrodynamics.
@@ 1 Introduction
Based on the original insights of Hohenberg and Kohn, who first proved that the ground-state electron density uniquely determines the energy of a many-electron system [66], density functional theory (DFT) has emerged as a standard tool in computational materials science. DFT’s success can be attributed to its ability to balance accuracy and computational cost [137]. At the heart of this balance lies the exchange-correlation (XC) functional, which approximates many-body electron interactions. This functional operates on a foundation that facilitates a ladder of approximations. The choice of XC approximation is critical, as it largely determines the trade-off between accuracy and computational expense in DFT simulations [27].
@@
The bottom rung in the ladder of functional approximations is the local density approximation (LDA), which approximates the exchange correlation energy based on a uniform electron gas model [92]. Advancements led to the formulation of generalized gradient approximations (GGA), which extend the LDA approach by incorporating gradient corrections to the electron density, enhancing the accuracy of DFT calculations. Several GGA functionals have been developed, as there are multiple ways to characterize the gradient of the electron density [92, 23, 93]. The two most common GGA functionals are the Perdew-Burke-Ernzerhof generalized gradient approximation (PBE) [92] and the modified Perdew-Burke-Ernzerhof generalized gradient approximation (PBEsol) [93]. PBE includes the gradient of the electron density in the exchange correlation function and improves upon LDA by capturing some non-local effects, but there remain limitations, particularly for systems with strongly localized electrons (e.g., transition metal complexes [148]). The third most common functional is meta-generalized gradient approximation (meta-GGA), which augments the functional dependence by including higher-order terms like the second derivative of the electron density [75]. In general, meta-GGAs (e.g., SCAN) offer improvements over GGAs in terms of accuracy for a variety of properties, including thermochemistry, and provide reliable predictions of structures and energies for a wide range of molecular and material bonding types [123]. Another noteworthy functional is the HSE (Heyd-Scuseria-Ernzerhof) hybrid functional which combines the traditional DFT exchange-correlation functional with a fraction of Hartree-Fock (HF) exchange [50] via a mixing parameter that determines the contribution of HF exchange to the total exchange-correlation energy. However, this functional requires greater computational resources than the simple PBE functional.
@@
Given these approximations, deviations between experimental and theoretical predictions are expected. It is well known that the LDA under-predicts the bandgap in silicon by about 50 % [94, 59, 151]. However, even with this significant discrepancy, accurate DFT-based predictions of the thermal conductivity of silicon have been obtained [73, 91, 83]. For instance, Jain and McGaughey showed that the choice of pseudopotential and XC functional in thermal conductivity predictions can impact accuracy by up to 17% [62].
Going beyond thermal conductivity, recent work has built upon these first principles calculations to investigate phonon hydrodynamics, a regime of thermal transport that lies between the diffusive and ballistic regimes [39]. One of the signatures of phonon hydrodynamics is the“second sound”, where the temperature propagates as a wave. Early observations of second sound were obtained for solid He4  [4], solid He3  [3] and NaF  [61]. Subsequent observations of second sound are scarce and limited to bismuth [85] and graphite  [57, 28]. Beyond these experimental observations of phonon hydrodynamics, there is a paucity of other experimental work. Rogers  [108] identified a narrow intermediate-temperature window in $^{7}\mathrm{LiF}$, where normal processes shape the propagated heat pulse, but the experimental evidence for phonon hydrodynamics remains inconclusive. Jelinek et al  [63] measured the thermal conductivity of $^{6}\mathrm{LiH}$ with 0.7% porosity from 6 to 399 K and observed a pronounced peak at approximately 50–70 K. Guyer and Krumhansl  [43] argued that isotope scattering in the then-available LiF crystals makes Poiseuille-type hydrodynamic flow unlikely, but second sound could still be observed at around 20 K in 2 mm sample radius, high-purity samples. Slack  [117] demonstrated that thermal conductivity in pure KCl at low temperatures is limited by boundary and isotope scattering, but no statement about the possibility of phonon hydrodynamics is made. A similar experimental and theoretical analysis was done for NaCl  [19]. To date, DFT has not been used to study phonon hydrodynamics in these materials. The objective of this work is to examine the impact of the choice of XC functionals (PBE, PBEsol, LDA) on the prediction of phonon hydrodynamics in fluorides, chlorides, and hydrides.
@@ 2 Theory
Based on approximate solutions to the phonon Boltzmann Transport Equation (BTE), a condition for the existence of phonon hydrodynamics was proposed by Guyer and Krumhansl, henceforth referred to as Guyer’s criterion [43]. This condition states that phonon hydrodynamics can be observed if the average boundary scattering rate $\langle\tau_{B}^{-1}\rangle$ is greater than the average Umklapp scattering rate $\langle\tau_{U}^{-1}\rangle$ but less than the average Normal scattering rate $\langle\tau_{N}^{-1}\rangle$ at a given temperature and for a characteristic length [43]:
$\langle\tau_{N}^{-1}\rangle\;>\;\langle\tau_{B}^{-1}\rangle\;>\;\langle\tau_{U}^{-1}\rangle,$ (1)
Here, $N$ denotes Normal (momentum-conserving) phonon–phonon scattering, $U$ denotes Umklapp (momentum-relaxing) phonon–phonon scattering, and $B$ denotes boundary scattering. In other words, phonon hydrodynamics can be observed if Normal scattering dominates over boundary scattering and Umklapp scattering. The average resistive scattering rate $\langle\tau_{R}^{-1}\rangle$ (sum of Umklapp and isotope scattering rates) can also be used instead of the average Umklapp scattering rate:
$\langle\tau_{N}^{-1}\rangle\;>\;\langle\tau_{B}^{-1}\rangle\;>\;\langle\tau_{R}^{-1}\rangle,$ (2)
@@
$\bigl\langle\tau^{-1}\bigr\rangle=\frac{\sum_{i=1}^{N}\tau_{i}^{-1}\,C_{i}}{\sum_{i=1}^{N}C_{i}}=\frac{\sum_{i=1}^{N}\tau_{i}^{-1}\,C_{i}}{C},$ (3)
where $C_{i}$ and $C$ are the heat capacity for a given phonon mode $i$ and the total heat capacity of all phonon modes and $N$ is the number of q-points in the discretized Brillouin zone, respectively. Other expressions for average scattering rates may be selected, but here we follow previous studies [22, 38, 39]. Extending our earlier study on NaF, LiF, LiH, and NaH [1, 56], we expand the analysis to a broader set of eight rock-salt compounds by including four additional materials and providing a more detailed, systematic comparison of XC functional effects on phonon transport and hydrodynamic windows. Details of the DFT and BTE calculations can be found in Appendix A of the supplementary material.
@@ 3.1 General Properties
All eight compounds studied in this work (NaF, LiF, KF, NaCl, KCl, LiH, NaH, and KH) crystallize in the rock-salt structure (face-centered cubic, space group $Fm\overline{3}m$) [54]. Each cation and anion is octahedrally coordinated by six neighbors. NaF, LiF, KF, KCl, LiH, NaH, and KH all adopt the NaCl-type lattice with similar octahedral frameworks, as shown in figure  1. A comprehensive analysis, given in table 1, of the structural, dielectric, vibrational, and electronic properties highlights the relative performance of the PBE, PBEsol, and LDA functionals. The calculated results are systematically compared against available experimental and theoretical data. For fluorides and chlorides, PBE slightly overestimates and LDA underestimates the lattice constant, while PBEsol is the closest to experiments, often deviating by less than 1% from experimental values. As GGA includes density-gradient corrections that reduce LDA’s systematic overbinding, it typically yields more accurate equilibrium volumes and lattice constants for real solids. For instance, PBEsol lattice constants for KF (5.31 Å) and KCl (6.24 Å) are in excellent agreement with their experimental measurement results of 5.31 Å and 6.25 Å [99]. Conversely, for the hydrides, the PBE functional tends to yield better results, as seen with LiH, where the PBE value (4.01 Å) is the closest to the experimental value of 4.08 Å [102]. The lattice constant increases monotonically from Li to Na to K for each anion, corresponding to the increasing ionic radii.
@@
The dielectric-constant tensor and Born effective charges are calculated based on the density functional perturbation theory (DFPT) to capture the longitudinal and transverse optical phonon (LO-TO) splitting at the gamma point [71]. For fluorides, chlorides and hydrides (NaF, LiF, LiH, etc.), the underlying Bravais lattice is cubic; hence, there is only one independent component in the Born effective charge tensor. In all cases, Born effective charges are nearly ionic ($\lvert Z^{\ast}\rvert\approx 1$), exemplified by NaCl (1.11, 1.09, 1.071.11,\,1.09,\,1.07 vs. $1.099$ [113]) and LiH (1.021.021.031.03 vs. $1.11$ [10]). For non-magnetic, transparent solids, a larger electronic polarizability leads to a larger refractive index ($n$) according to the Lorentz–Lorenz relation: $(n^{2}-1)/(n^{2}+2)=N\alpha/(3\varepsilon_{0})$, and in the optical limit with $\mu_{r}\approx 1$ the high-frequency dielectric constant satisfies $\varepsilon_{\infty}\approx n^{2}$ [41]. Across anions at fixed cation, the electronic polarizability increases in the sequence $\mathrm{F^{-}}<\mathrm{Cl^{-}}<\mathrm{H^{-}}$, so the high-frequency dielectric $\varepsilon_{\infty}$ constant and thus the refractive index ($n\approx\sqrt{\varepsilon_{\infty}}$ ) rises accordingly. The calculated high-frequency dielectric constants follow chemical intuition: fluorides are the least polarizable ($\varepsilon_{\infty}$ = 1.81.82.12.1), chlorides are higher ($\varepsilon_{\infty}$ = 2.32.32.72.7), and hydrides are the largest ($\varepsilon_{\infty}$ = 2.72.74.84.8).
@@
A consistent weakness remains in the prediction of the electronic band gap ($E_{g}$) using these functionals. For example, the calculated values for KF ($\sim 6\,\mathrm{eV}$) and KCl ($\sim 5\,\mathrm{eV}$) are significantly lower than their respective experimental measurements of $10.90\,\mathrm{eV}$ and $8.70\,\mathrm{eV}$ [90]. It was previously found that other functionals, such as the modified Becke–Johnson (mBJ) exchange potential, can be used to obtain a more accurate estimate of the band gap [136, 20, 78]. The electronic band structures of fluoride, chloride and hydride materials can be found in Appendix B of the supplementary material, where the Fermi energy ($E_{f}$) is shifted to zero eV.
@@
The dispersion curves for the eight fluoride, chloride, and hydrides, seen in figure  2, are qualitatively consistent across PBE, PBEsol, and LDA and align with the experimental measurement results. The phonon frequency values of LO and TO modes at the gamma point are also given in table 1. LDA leads to slightly higher phonon frequencies compared to PBE and PBEsol. Nevertheless, the predictions from all three functionals are in good agreement with experimental measurements along high-symmetry lines. A clear hierarchy emerges from the band shapes. Within each anion family, replacing Li with Na and then K progressively softens both acoustic and optical branches and lowers sound velocities. Across a fixed cation, moving from chloride to fluoride to hydride pushes the optical manifold upward and opens a pronounced acoustic–optical gap. The hydrides (LiH, NaH, KH) show the widest gap and the flattest high-frequency optic branches, reflecting the large mass contrast and strong long-range Coulomb coupling [128, 30]. The large LO-TO splitting frequency values are due to strong polar phonon frequencies (creation of an oscillating electric field inside a crystal caused by phonons due to local charge polarization) [6]. The $\Gamma$-point LO–TO separations quantify these patterns. Using the tabulated PBE values, the hydrides have the largest splittings, with $\Delta_{\text{LO--TO}}$ of $14.87\,\mathrm{THz}$ for LiH, $11.78\,\mathrm{THz}$ for NaH, and $9.85\,\mathrm{THz}$ for KH. Fluorides are intermediate, with $10.54\,\mathrm{THz}$ for LiF, $4.89\,\mathrm{THz}$ for NaF, and $4.07\,\mathrm{THz}$ for KF. Chlorides are the smallest, with $2.87\,\mathrm{THz}$ for NaCl and $2.12\,\mathrm{THz}$ for KCl. The PDOS curves reflect these trends.
@@
Figure 2: Calculated phonon dispersion and phonon density of states for: (a) NaF, black circles are the measured data from Ref. [18]. (b) LiF, black circles are the measured data from Ref. [29]. (c) KF, black circles are the measured data from Ref. [17]. (d) NaCl, black circles are the measured data from Ref. [111]. (e) KCl, black circles are the measured data from Ref. [25]. (f) LiH, black circles are the measured data from Ref. [139]. (g) NaH. (h) KH. Calculations were done with PBE (black), PBEsol (red), and LDA (blue).
@@
Table 1: Lattice constant ($a$), high-frequency dielectric constant ($\varepsilon_{\infty}$), Born effective charge ($Z^{*}$), phonon frequencies at $\Gamma$ ($\omega$), and band gap ($E_{g}$) of calculated (PBE, PBEsol, LDA), other theoretical data, and experimental data for NaF, LiF, KF, NaCl, KCl, LiH, NaH and KH.
@@
$Z^{*}$ 1.02 1.01 0.99 1.02e, 0.96i 1.02j
$\omega_{\mathrm{LO}}$ (THz) 11.67 12.01 12.81 11.90e 12.65k
$\omega_{\mathrm{TO}}$ (THz) 6.78 7.24 8.32 7.70c, 6.95e 7.86h, 7.51k
$E_{g}$ (eV) 6.02 6.11 6.34 12.00a, 6.12 (GGA)e, 7.11 (EV)e, 11.69 (mBJ)e, 5.94l 11.50m
@@
$Z^{*}$ 1.05 1.05 1.04 0.99i, 1.03o, 1.06p
$\omega_{\mathrm{LO}}$ (THz) 18.63 19.18 20.09 19.89o, 19.92p 19.70t
$\omega_{\mathrm{TO}}$ (THz) 8.09 9.06 10.28 9.00c, 10.61o, 10.29p 9.54h, 9.15t
$E_{g}$ (eV) 8.82 9.04 9.64 9.11l, 8.73p, 7.64 (LDA)u, 9.15 (PBE)u, 8.70v 13.60w, 14.20x
@@
$\omega_{\mathrm{LO}}$ (THz) 9.05 9.45 10.18 9.55aa 9.55y
$\omega_{\mathrm{TO}}$ (THz) 4.98 5.42 6.33 5.80c, 5.89aa 6.05h, 5.57y
$E_{g}$ (eV) 5.96 6.15 6.36 6.14l 10.90ab
@@
$a$ (Å) 5.70 5.61 5.47 5.80c, 5.70d, 5.70 (FP -LAPW)e, 5.70 (PP)e, 5.52ac, 5.70ad, 5.47 (LDA)ae, 5.70 (GGA)ae 5.57c, 5.60f, 5.60ac, 5.60ae
$\varepsilon_{\infty}$ 2.48 2.57 2.67 2.47e, 2.55af 2.35h
$Z^{*}$ 1.11 1.09 1.07 1.11e, 0.99i, 1.10af 1.10i
$\omega_{\mathrm{LO}}$ (THz) 7.34 7.60 8.07 7.46e 7.93ag
$\omega_{\mathrm{TO}}$ (THz) 4.47 4.93 5.63 4.70c, 4.62e 5.34h, 5.17ag
$E_{g}$ (eV) 4.98 4.99 5.04 5.00l, 5.00 (GGA)e, 6.07 (EV)e, 8.41 (mBJ)e 8.75ah, 9.00ai
@@
$a$ (Å) 6.38 6.24 6.07 6.57c, 6.40d, 6.30ac 6.25f, 6.24ac
$\varepsilon_{\infty}$ 2.28 2.38 2.51 2.37af 2.20h, 2.15aj
$Z^{*}$ 1.13 1.12 1.11 1.13af
$\omega_{\mathrm{LO}}$ (THz) 5.89 6.16 6.66 6.34aj
$\omega_{\mathrm{TO}}$ (THz) 3.77 4.15 4.85 4.10c, 4.41aj 4.53h, 4.36aj
$E_{g}$ (eV) 5.05 5.11 5.12 5.32l, 5.50ak 8.70ab
@@
$Z^{*}$ 1.02 1.02 1.03 1.05i, 1.02am, 1.03ar 1.11ar, 0.99as
$\omega_{\mathrm{LO}}$ (THz) 32.43 32.37 32.96 28.30ar 32.21an, 33.50ar, 33.60as, 33.58at
$\omega_{\mathrm{TO}}$ (THz) 17.56 18.52 19.41 18.10ar, 17.86au 18.14an, 18.40ar, 17.76as, 17.70at
$E_{g}$ (eV) 2.98 2.73 2.54 3.00 (GGA)al, 2.61 (LDA)al, 4.75 (GW)al, 2.53au 5.00av, 4.92aw
@@
$Z^{*}$ 0.97 0.96 0.95 0.97am, 0.96ax, 0.99 (GGA)ba, 0.97 (LDA)ba
$\omega_{\mathrm{LO}}$ (THz) 26.75 26.65 27.52 27.06am
$\omega_{\mathrm{TO}}$ (THz) 14.97 15.44 16.83 16.60 (CN)az, 13.90 (QH)az 15.00az
$E_{g}$ (eV) 3.76 3.61 3.55 3.79 (GGA)al, 3.42 (LDA)al, 5.87 (GW)al, 4.90ax, 5.68 (GW)bc, 3.39 (LDA)bc